A Clinical Psychologist wants to know what color is the most relaxing, to design new “calm” waiting rooms for dentists. Three different colors are viewed by participants and rated on a scale (0 – 10) to quantify the level of relaxation caused by the color. The n value and mean for each color are shown below, higher scores mean greater relaxation. Complete the source table below and answer the accompanying questions. Red Green Blue n = 9 n = 9 n = 9 M = 1.9 M = 3.2 M = 4.3 Source SS df MS Fcalculated Fcritical p level Between Treatments ___28__ ____ _____ ______ _____ p __ .05 Within Treatments __110_ ____ _____ Total _______ _____ What was the dependent variable ( D.V. ) in this study? __________________ What was the independent variable (I.V.) in this study? ___________________ Was there a significant effect? ______________________________________ Write an APA formatted summary of this study below.
Answer #1

Source                                   SS            df              MS        Fcalculated       Fcritical         p level         

Between Treatments     ___28__     __2__       __14___ ______3.057 ____3.40_       p __ .05     

Within Treatments         __110_       __24__       __4.58___         

Total _______ _____

df Between Treatments = Number of groups - 1 = 3 - 1 = 2

df Within Treatments = Number of observations - Number of groups = 9 * 3 - 3 = 24

MS Between Treatments = SS Between Treatments / df Between Treatments = 28 / 2 = 14

MS Within Treatments = SS Within Treatments / df Within Treatments = 110 / 24 = 4.58

Fcalculated = MS Between Treatments / MS Within Treatments = 14 / 4.58 = 3.057

Fcritical  for df = 2, 24 and 0.05 significance level is 3.40

  1. What was the dependent variable ( D.V. ) in this study is level of relaxation. __________________
  2. What was the independent variable (I.V.) in this study is color. ___________________
  3. Was there a significant effect.  Since the Fcalculated < Fcritical , there was not a significant effect.
  4. Write an APA formatted summary of this study below.  A one-way between subjects ANOVA was conducted to compare the effect of color on level of relaxation with Red, Green and Blue conditions. There was not a significant) effect of color on level of relaxation at the p<.05 level for the three conditions [F(2, 24) = 3.057, p = 0.0657